OpenStudy (anonymous):

Determine the slope of (a) every line that is parallel to the given line and (b) every line 2y-5x+4=0

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

Bazinga! All in my head! :D

OpenStudy (anonymous):

but here is how I got that: 2y-5x+4=0 2y=5x-4 y = (5/2)x + 2 Because it is in the form y=mx+b, m (the slope) is 5/2. Parallel lines will have the same slope.

OpenStudy (anonymous):

That is what parallel is.

OpenStudy (anonymous):

That is awesome. How do I determine the perpendicular line?

OpenStudy (anonymous):

Wow, that last response posted way too many times; i apologize, that was not me i promise.